OpenStudy (anonymous):

Please Help! Best Response! B = 36°, a = 41, c = 20 b ≈ 27.5, C ≈ 115.5, A ≈ 28.5 b ≈ 45.3, C ≈ 25.5, A ≈ 118.5 b ≈ 45.3, C ≈ 29.5, A ≈ 114.5 b ≈ 27.5, C ≈ 25.5, A ≈ 118.5

hartnn (hartnn):

did you try this? if yes, where r u stuck? do you know sine law ?

OpenStudy (anonymous):

I do know applying it to problems like this one

hartnn (hartnn):

ok, let me tell you both the laws which you can use here : SINE LAW: \(\dfrac{\sin A}{a}=\dfrac{\sin B}{b}=\dfrac{\sin C}{c}\) COS LAW : \(b^2=a^2+c^2+2ac \cos B\) do you know whic law/formula you can use first ?

hartnn (hartnn):

sorry its, \(b^2=a^2+c^2-2ac \cos B\)

OpenStudy (anonymous):

so itd b b^2 = 41^2 + 20^2 - 2(41) cos 36?

hartnn (hartnn):

you missed 'c' =20 b^2 = 41^2 + 20^2 - 2(41) (20) cos 36 otherwise it was correct, so you can find 'b' directly from here! can u?

OpenStudy (anonymous):

Is it 755?

hartnn (hartnn):

thats is b^2, so b =.... ?

OpenStudy (anonymous):

27.4

OpenStudy (anonymous):

so how do i find C?

hartnn (hartnn):

yes, approx its b=27.5 good :)

hartnn (hartnn):

now you can use sine law

hartnn (hartnn):

you know b, sin B and c so, you can find sin C \(\dfrac{\sin B}{b}=\dfrac{\sin C}{c}\)

OpenStudy (anonymous):

I got 25.5 :)

hartnn (hartnn):

yes! now again using sine law, finish it by findind A also.

OpenStudy (anonymous):

Okay thank you so much!, you helped me a lot! :)

hartnn (hartnn):

welcome ^_^